# Reactive Diffusion Policy: Slow-Fast Visual-Tactile Policy Learning for Contact-Rich Manipulation

### v2

@@ Abstract
Humans can accomplish complex contact-rich tasks using vision and touch, with highly reactive capabilities such as quick adjustments to environmental changes and adaptive control of contact forces; however, this remains challenging for robots. Existing visual imitation learning (IL) approaches rely on action chunking to model complex behaviors, which lacks the ability to respond instantly to real-time tactile feedback during the chunk execution. Furthermore, most teleoperation systems struggle to provide fine-grained tactile / force feedback, which limits the range of tasks that can be performed. To address these challenges, we introduce TactAR, a low-cost teleoperation system that provides real-time tactile feedback through Augmented Reality (AR), along with Reactive Diffusion Policy (RDP), a novel slow-fast visual-tactile imitation learning algorithm for learning contact-rich manipulation skills. RDP employs a two-level hierarchy: (1) a slow latent diffusion policy for predicting high-level action chunks in latent space at low frequency, (2) a fast asymmetric tokenizer for closed-loop tactile feedback control at high frequency. This design enables both complex trajectory modeling and quick reactive behavior within a unified framework. Through extensive evaluation across three challenging contact-rich tasks, RDP significantly improves performance compared to state-of-the-art visual IL baselines through rapid response to tactile / force feedback. Furthermore, experiments show that RDP is applicable across different tactile / force sensors. Code and videos are available on reactive-diffusion-policy.github.io .
@@ I Introduction
To leverage the high-quality visual tactile data collected by the TactAR system, we propose an imitation learning algorithm called Reactive Diffusion Policy (RDP) (Fig. 1 right). Inspired by human control strategies in contact-rich tasks, we adopt a “slow-fast” hierarchical control pipeline. We use a slow network (latent diffusion policy, LDP) to predict a high-level action chunk in latent space at low frequency (1-2 Hz), which is analogous to the predictive action. Then we use a fast network (Asymmetric Tokenizer, AT) to finetune the latent action chunk based on high-frequency tactile feedback (20-30 Hz), which is analogous to the closed-loop finetuning. This hierarchical structure allows the slow network to preserve its ability to model complex and non-Markovian actions via the diffusion model [10] and action chunking, while the fast network achieves closed-loop control with real-time tactile feedback for precise force control and quick response.
We have conducted experiments on three challenging contact-rich tasks, which can evaluate the model performance on the following aspects: (1) precision, (2) precise and adaptive force control, (3) fast response under disturbances, and (4) bimanual coordination. Real-world experiments have shown that our Reactive Diffusion Policy algorithm can model complex actions while maintaining very fast reactive behavior, achieving a significant performance improvement ($>35\%$) in three tasks compared to state-of-the-art IL baselines. This also validates the data quality of our TactAR teleoperation system. Furthermore, we conduct cross-sensor experiments demonstrating that RDP is applicable across various tactile / force sensors (GelSight Mini[20], MCTac [48] and built-in joint torque sensors [19]). The code of the TactAR teleoperation system and Reactive Diffusion Policy algorithm are available at reactive-diffusion-policy.github.io .
@@ III Teleoperation System: TactAR
  • Low-cost. The system do not need specialized or isomorphic hardwares for haptic feedback. It only needs a consumer-level VR headeset (Meta Quest3) for teleoperation, which costs only \$$\$500$.
@@ V-A5 Implementation Details
The Diffusion Policy and our slow policy (LDP) predict open-loop 12 FPS action sequences for each action chunk. The low-frequency observation of LDP includes both visual inputs and tactile / force inputs. The fast policy (AT) takes tactile / force observations at 24 FPS and update new action predictions at 24 FPS. Note that we use 24 FPS because we are constrained by the frame rate limitation of GelSight [20], which is 25 FPS. Our RDP algorithm can also be applied to higher frequency tactile / force signals in theory. Please see Appendix -D, -F and -H for more details on data collection, the inference process and the hyperparameters.
@@ V-B Results
We also found that Diffusion Policy (DP) incorporates tactile embedding (DP w. tactile emb.) performs similarly compared to DP with purely visual inputs in the three tasks (see Tab II, Tab III and Tab IV). However, despite similar performance, these two DP baselines exhibit different failure modes. We observe that DP with pure visual input frequently predicts inaccurate trajectories and results in large contact forces (e.g., failure case 2 in Fig. 8 (b)), necessitating the human evaluators to move their hands to prevent sensor damage. In contrast, DP with tactile embedding rarely brings large contact forces. It may get stuck when making contact with the object (e.g., failure case 2 in Fig. 8 (a) & failure case in Fig. 8 (c)). This may be because it learns some reactive behavior from the data (e.g., slightly move up when making contact; move down when loose contact). However, since it executes action chunks in an open-loop manner, it lacks the capability for finer adjustments. Consequently, it might quickly transition to a different contact state and repeatedly switch between various contact states. In conclusion, we need a better way to incorporate tactile signals into the policy.
@@
RDP is also applicable for force / torque sensors. Tab. II, Tab. III and Tab. IV have shown that RDP with simple force input without any additional design achieves the best results in all three tasks. Despite the significant noise associated with the force sensor during rapid robot movements (as evidenced in the supplementary video’s content for TactAR), the RDP algorithm successfully identifies useful patterns from the noisy data. We hypothesize that the force sensor may yield better results due to its lower latency and reduced dimensionality compared to optical tactile sensors, which potentially facilitates easier learning by the network.
@@
Slow-fast hierarchy, relative trajectory and latency matching are essential for RDP performance (Q5). As shown in Fig. 5, there are two ways to increase the closed-loop control frequency without our slow-fast hierarchy: (1) reducing action chunk size. (2) using temporal ensemble. However, experiments in Tab. V have proved that these two options both have significant side effects. We can see from Tab. V that when the action chunk size is reduced from 8 to 2, the DP baseline tends to get stuck before grasping (failure case 4 in Fig. 8 (b)), which makes the grasp success rate drop from $100\%$ to $20\%$. Policy with small chunk size is very sensitive to non-Markovion behaviors (e.g., pauses in the air) commonly found in human demonstration data, so we can not simply reduce chunk size.
@@
Tactile / force feedback in TactAR improves data quality in contact-rich tasks by improving the stability of contact forces (Q6). We have conducted a user study on how tactile / force feedback in TactAR helps the data collection process. We invited 10 users with different levels of experience in VR teleoperation or Imitation Learning (IL). Please see Appendix -G for more details of the user study. We perform $10\times 10\times 2=200$ trials in total for quantitative analysis. As shown in Fig. 11, most of the users ($\geq 70\%$) found that tactile / force AR feedback is very helpful in data collection, regardless of whether they were acting as the teleoperator or the person holding the cucumber. In Fig. 12, we can see that tactile / force feedback in TactAR can greatly improve the data quality from both the normalized peeling length ($0.72\rightarrow 0.91$) and the ratio of stable contact force ($0.58\rightarrow 0.87$).
@@
Data with higher stability of contact forces will help the model to discover useful patterns more easily (Q7). We have collected the same number of demonstrations (60) for Peeling task with traditional VR teleoperation without tactile / force feedback and trained RDP (force) with these data. The results in Fig. 14 show that data quality has a large influence on policy performance (the score decreases by more than $30\%$). We observe that policies trained with low-quality data exhibited more unstable performance, such as more unstable force during peeling and a higher likelihood of breaking halfway through. A possible explanation is that the Fast Policy in RDP is designed to identify associations between tactile / force signals and trajectories from the data and learn reactive behavior. When contact forces in the data are highly unstable, the Fast Policy struggles to identify reasonable associations, which reduces performance.

### v3

@@ Abstract
Humans can accomplish complex contact-rich tasks using vision and touch, with highly reactive capabilities such as fast response to external changes and adaptive control of contact forces; however, this remains challenging for robots. Existing visual imitation learning (IL) approaches rely on action chunking to model complex behaviors, which lacks the ability to respond instantly to real-time tactile feedback during the chunk execution. Furthermore, most teleoperation systems struggle to provide fine-grained tactile / force feedback, which limits the range of tasks that can be performed. To address these challenges, we introduce TactAR, a low-cost teleoperation system that provides real-time tactile feedback through Augmented Reality (AR), along with Reactive Diffusion Policy (RDP), a novel slow-fast visual-tactile imitation learning algorithm for learning contact-rich manipulation skills. RDP employs a two-level hierarchy: (1) a slow latent diffusion policy for predicting high-level action chunks in latent space at low frequency, (2) a fast asymmetric tokenizer for closed-loop tactile feedback control at high frequency. This design enables both complex trajectory modeling and quick reactive behavior within a unified framework. Through extensive evaluation across three challenging contact-rich tasks, RDP significantly improves performance compared to state-of-the-art visual IL baselines. Furthermore, experiments show that RDP is applicable across different tactile / force sensors. Code and videos are available on reactive-diffusion-policy.github.io.
@@ I Introduction
To leverage the high-quality visual tactile data collected by the TactAR system, we propose an imitation learning algorithm called Reactive Diffusion Policy (RDP) (Fig. 1 right). Inspired by human control strategies in contact-rich tasks, we adopt a “slow-fast” hierarchical control pipeline. We use a slow network (latent diffusion policy, LDP) to act like a neural planner and predict a high-level action chunk in latent space at low frequency (1-2 Hz), which is analogous to the predictive action. Then we use a fast network (Asymmetric Tokenizer, AT) to act like a learnable impedance controller and finetune the latent action chunk based on high-frequency tactile feedback (20-30 Hz), which is analogous to the closed-loop finetuning. This hierarchical structure allows the slow network to preserve its ability to model complex and non-Markovian actions via the diffusion model [10] and action chunking, while the fast network achieves closed-loop control with real-time tactile feedback for precise force control and quick response.
We have conducted experiments on three challenging contact-rich tasks, which can evaluate the model performance on the following aspects: (1) precision, (2) precise and adaptive force control, (3) fast response under disturbances, and (4) bimanual coordination. Real-world experiments have shown that our Reactive Diffusion Policy algorithm can model complex actions while maintaining very fast reactive behavior, achieving a significant performance improvement ($>35\%$) in three tasks compared to state-of-the-art IL baselines. We have also conducted user studies and quantitative analyses to validate the high quality of the data collected by TactAR teleoperation system. Furthermore, we conduct cross-sensor experiments demonstrating that RDP is applicable across various tactile / force sensors (GelSight Mini[20], MCTac [48] and built-in joint torque sensors [19]). The code of the TactAR teleoperation system and Reactive Diffusion Policy algorithm are available on reactive-diffusion-policy.github.io.
@@ III Teleoperation System: TactAR
  • Low-cost. The system do not need specialized or isomorphic hardwares for haptic feedback. It only needs a consumer-level VR headset (Meta Quest3) for teleoperation, which costs only \$$\$500$.
@@ V-A5 Implementation Details
The Diffusion Policy and our slow policy (LDP) predict open-loop 12 FPS action sequences for each action chunk. The low-frequency observation of LDP includes both visual inputs and tactile / force inputs. The fast policy (AT) takes tactile / force observations at 24 FPS and update new action predictions at 24 FPS. Note that we use 24 FPS because we are constrained by the frame rate limitation of GelSight [20], which is 25 FPS. Our RDP algorithm can also be applied to higher frequency tactile / force signals in theory. Please see Appendix -D, -F and -I for more details on data collection, the inference process and the hyperparameters.
@@ V-B Results
We also find that Diffusion Policy (DP) incorporates tactile embedding (DP w. tactile emb.) performs similarly compared to DP with purely visual inputs in the three tasks (see Tab II, Tab III and Tab IV). However, despite similar performance, these two DP baselines exhibit different failure modes. We observe that DP with pure visual input frequently predicts inaccurate trajectories and results in large contact forces (e.g., failure case 2 in Fig. 8 (b)), necessitating the human evaluators to move their hands to prevent sensor damage. In contrast, DP with tactile embedding rarely brings large contact forces. It may get stuck when making contact with the object (e.g., failure case 2 in Fig. 8 (a) & failure case in Fig. 8 (c)). This may be because it learns some reactive behavior from the data (e.g., slightly move up when making contact; move down when loose contact). However, since it executes action chunks in an open-loop manner, it lacks the capability for finer adjustments. Consequently, it might quickly transition to a different contact state and repeatedly switch between various contact states. In conclusion, we need a better way to incorporate tactile signals into the policy.
@@
RDP is also applicable for force / torque sensors. Tab. II, Tab. III and Tab. IV have shown that RDP with simple force input without any additional design achieves the best results in all three tasks. Despite the significant noise associated with the force sensor during rapid robot movements (as evidenced in the TactAR part of the supplementary video), the RDP algorithm successfully identifies useful patterns from the noisy data. We hypothesize that the force sensor may yield better results due to its lower latency and reduced dimensionality compared to optical tactile sensors, which potentially facilitates easier learning by the network.
@@
Slow-fast hierarchy, relative trajectory and latency matching are essential for RDP performance (Q5). As shown in Fig. 5, there are two ways to increase the closed-loop control frequency without our slow-fast hierarchy: (1) reducing action chunk size. (2) using temporal ensemble. However, experiments in Tab. V have proved that these two options both have significant side effects. We can see from Tab. V that when the action chunk size is reduced from 8 to 2, the DP baseline tends to get stuck before grasping (failure case 4 in Fig. 8 (b)), which makes the grasp success rate drop from $100\%$ to $20\%$. Policy with small chunk size is very sensitive to non-Markovian behaviors (e.g., pauses in the air) commonly found in human demonstration data, so we can not simply reduce chunk size.
@@
Please see Appendix -G for more ablation results.
Tactile / force feedback in TactAR improves data quality in contact-rich tasks by improving the stability of contact forces (Q6). We have conducted a user study on how tactile / force feedback in TactAR helps the data collection process. We invited 10 users with different levels of experience in VR teleoperation or Imitation Learning (IL). Please see Appendix -H for more details of the user study. We perform $10\times 10\times 2=200$ trials in total for quantitative analysis. As shown in Fig. 11, most of the users ($\geq 70\%$) found that tactile / force AR feedback is very helpful in data collection, regardless of whether they were acting as the teleoperator or the person holding the cucumber. In Fig. 12, we can see that tactile / force feedback in TactAR can greatly improve the data quality from both the normalized peeling length ($0.72\rightarrow 0.91$) and the ratio of stable contact force ($0.58\rightarrow 0.87$).
@@
High-quality data will help the model to discover useful patterns more easily (Q7). We have collected the same number of demonstrations (60) for Peeling task with traditional VR teleoperation without tactile / force feedback and trained RDP (force) with these data. The results in Fig. 14 show that data quality has a large influence on policy performance (the score decreases by more than $30\%$). We observe that policies trained with low-quality data exhibited more unstable performance, such as more unstable force during peeling and a higher likelihood of breaking halfway through. A possible explanation is that the Fast Policy in RDP is designed to identify associations between tactile / force signals and trajectories from the data and learn reactive behavior. When contact forces in the data are highly unstable, the Fast Policy struggles to identify reasonable associations, which reduces performance.
@@ Acknowledgments
This work is supported by the Shanghai Commitee of Science and Technology, China(Grant No.24511103200) by the National Key Research, Development Project of China (No. 2022ZD0160102), XPLORER PRIZE grants and Tsinghua Dushi program.